OpenStudy (queenb__):

The data table represents the distance between a well-known lighthouse and a cruise ship over time. The cruise ship is travelling at uniform speed. Which equation represents the distance (y) from the lighthouse, based on the number of hours (x)? Number Distance from Lighthouse of Hours (in oceanic miles) 2 53 4 95.5 6 138 8 180.5 10 223 12 265.5 14 308 16 350.5

OpenStudy (queenb__):

y = 42.5x + 10.5 y = 10.5x + 32 y = 21.25x + 10.5 y = 12.25x + 28.5 y = 12.5x + 10.5

563blackghost (563blackghost):

Find the slope between two points. `(hours,miles)`. So we will use `(2,53) and (4,95.5)` \(\huge\bf{\frac{rise}{run} \rightarrow \frac{y_{2}-y_{1}}{x_{2}-x_{1}} \rightarrow \frac{95.5 - 53}{4-2}}\)

OpenStudy (queenb__):

y = 21.25x + 10.5 @563blackghost